OpenStudy (anonymous):

I need help with percentages. How do I solve: $37.50 per pack of 100. What is cost of 300 packs with a 15% discount?

OpenStudy (anonymous):

without discount 300 packs cost $37.50 * 300 = $11,250 15 % discount = $11,250 * 15 / 100 = $1687.50 so cost after discount = 11250 - 1687.50 = $9562.50

OpenStudy (anonymous):

Thank you so much for showing me how to get the answer.

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

basically, percentage is chaged to fraction by dividing by 100 5 % = 5/100 etc